OpenStudy (anonymous):

Order 1/3,1/2 ,3/4 and from least to greatest.

OpenStudy (raden):

1/3 = 3/12 1/2 = 6/12 3/4 = 9/12 now, just compare who is the greatest of the numerators

OpenStudy (anonymous):

First you want to make them all like fractions so 4/12, 6/12, 9/12 so the answer is \[\left( \frac{ 1 }{ 3 }\frac{ 1 }{ 2 }\frac{ 3 }{ 4 } \right)\]

OpenStudy (anonymous):

\[\frac{ 1 }{ 3 } \approx 0.33333333333333333\] \[\frac{ 1 }{ 2 } = 0.5 \] \[\frac{ 3 }{ 4 } = 0.75 \]